OpenStudy (anonymous):

What is the correct radical form of this expression? (32a^10b^5/2)^2/5

OpenStudy (anonymous):

@rishavraj

rishavraj (rishavraj):

is it \[(32a ^{10} \times 10b ^{\frac{ 5 }{ 2 }})^{\frac{ 2 }{ 5 }}\]

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

Wait let me write down the options.

rishavraj (rishavraj):

a bit mistake it should be \[2^2 a^4 b\]

rishavraj (rishavraj):

i simplified it...lol answer would be A.... \[\sqrt[a]{x} = x ^{\frac{ 1 }{ a }}\]

OpenStudy (anonymous):

Thank you.

rishavraj (rishavraj):

u r most welcome...:)